OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation @agent0smith please

OpenStudy (anonymous):

(x+5)^2/36-(y+1)/64=1

OpenStudy (agent0smith):

Find the center first: (-5, -1) vertices \[\Large (-5 \pm a, -1)\]where a^2 is the number under the x^2 term. b^2 is under the y term. foci \[\Large (-5 \pm c, -1)\] where c^2 = a^2+b^2

OpenStudy (anonymous):

Vertices: (-1, 1), (-1, -11); Foci: (-1, -15), (-1, 5)?

OpenStudy (ranga):

Is the (y+1) term squared? (y+1)^2 ? Otherwise this is a parabola :)

OpenStudy (anonymous):

yes @ranga

OpenStudy (agent0smith):

You seem to have mixed the x and y's up... see my post again.

OpenStudy (anonymous):

@agent0smith so would it be:

OpenStudy (anonymous):

Vertices: (1, -1), (-11, -1); Foci: (-15, -1), (5, -1)

OpenStudy (agent0smith):

Looks fine.